OpenStudy (anonymous):

if the difference between the simple interest and compound interest compounded annually on a certain sum for two years at 5% rate of interest is rupee 1,then the sum is

OpenStudy (kropot72):

Let the sum = P The simple interest on the sum for 2 years = 2 * 0.05P If the sum has interest compounded for 2 years, the total amount of the sum and the interest is given by \[A=P(1+0.05)^{2} .........(1)\] The compound interest on the sum for 2 years is therefore \[1.05^{2}P-P=P(1.05^{2}-1)=0.1025P\] The difference between the amounts of compound interest and simple interest is given by \[0.1025P-0.1P=0.0025P=1\ rupee\] Therefore the sum P is found from: \[P=\frac{1}{0.0025}\ rupees\]